OpenStudy (anonymous):

4 1/3*7 3/4= 403/12= 396 7/12 is this right

OpenStudy (anonymous):

yup, spot on

OpenStudy (anonymous):

and 6-2 4/5=is 16/5 which is 1 3/5 right

OpenStudy (anonymous):

not quite. 16/5 is right tho, but 1 3/5 is actually 8/5

OpenStudy (anonymous):

I suspect you switched the 1 and the 3 now that I think of it

OpenStudy (anonymous):

wait im confused

OpenStudy (anonymous):

16/5 is correct, but you seem to have made a mistake after that. 16/5 = (5+5+5+1)/5 = 3 1/5

OpenStudy (anonymous):